Q3 A binary vapor mixture contains 65.00% ethanol (1) in water (2) at 50 oC. Determine the pressure in which this vapor develops the first drop of liquid b) a) the liquid composition Use NRTL to calculate the activity coefficient a12 3437 cal/mol, a2 1022.9466 cal/mol and a12-0.2935 Note: To solve this exercise, you will need to assign a number to x and calculate for psat XY Psat y1.cal =-iriri sat Then, goalseek 이.cal-y,actual) to zero by changing xỉ.
Answer #1

We will be using iterative procedure for calculating Xi , NRTL eqns are given as;

2 G21 T21 G12T12 1 T2721

\ln \gamma _{1}=x_{2}^{2}\left [ \tau _{21}\left (\frac{G_{21}}{x_{1}+x_{2}G_{21}} \right )^{2}+\frac{G_{12}\tau _{12}}{(x_{2}+x_{1}G_{12})^{2}}\right ] ......... activity coefficient of component 1

\ln \gamma _{2}=x_{1}^{2}\left [ \tau _{12}\left (\frac{G_{12}}{x_{2}+x_{1}G_{12}} \right )^{2}+\frac{G_{21}\tau _{21}}{(x_{1}+x_{2}G_{21})^{2}}\right ]

G_{12}=\exp (-\alpha \tau _{12}) ............. G_{21}=\exp (-\alpha \tau _{21})

\tau _{21}=\frac{a_{21}}{RT}..........\tau _{12}=\frac{a_{12}}{RT}

Using all these formulas and given information   y_{1}=0.65..............y_{2}=0.35

Pressure at which first drop appears is dew point pressure;

first calculate vapou pressure of components at 50o C

\ln P^{Sat}(KPa)=A-\frac{B}{T+C}........... a,b,c ..parameters

for ethanol A = 16.8958 B = 3795.17 C = -42.232

for Water A = 16.3872 B = 3885.7 C = -42.98

Psat (ethanol)= 29.545 KPa Psat (water)= 12.405 KPa  

Intially Assume x1 = x2 = 0.5

\tau _{21}=1.594.........\tau _{12}=0.0176

G_{12}=0.995 ............. G_{21}=0.626

\gamma _{1}=1.2665....\gamma _{2}=1.4646

P^{Tot}=\sum \gamma _{i}x_{i}P_{i}^{Sat}\Rightarrow 27.79 KP a

y_{1}=0.67....... y_{2}=0.33

x_{i}=\frac{y_{i}P^{Tot}}{\gamma _{i}P_{i}^{Sat}}\Rightarrow x_{1}=0.497..and..x_{2}=0.503

Using these values again repeat the procedure;

\gamma _{1}=1.6368....\gamma _{2}=1.46

P^{Tot}=\sum \gamma _{i}x_{i}P_{i}^{Sat}\Rightarrow 33.144 KP a

y_{1}=0.72....... y_{2}=0.28

x_{i}=\frac{y_{i}P^{Tot}}{\gamma _{i}P_{i}^{Sat}}\Rightarrow x_{1}=0.493..and..x_{2}=0.507\

again iterating;

\gamma _{1}=1.2827....\gamma _{2}=1.4527

P^{Tot}=\sum \gamma _{i}x_{i}P_{i}^{Sat}\Rightarrow 27.81 KP a

y_{1}=0.672....... y_{2}=0.328

x_{i}=\frac{y_{i}P^{Tot}}{\gamma _{i}P_{i}^{Sat}}\Rightarrow x_{1}=0.493..and..x_{2}=0.507

Therefore ; the dew point pressure is

P^{Dew }= 27.81 KPa....................(a)

Liquid composition

Ethanol(x_{1}) = 0.493.......... Water(x_{2}) = 0.507......................(b)